OpenStudy (anonymous):

please help Classify the following system of equations: 9x − 3y = 12 2y = 6x − 8 Consistent and independent Inconsistent Consistent and dependent

OpenStudy (anonymous):

9x − 3y = 12 2y = 6x − 8 3x -y =4 6x -2y =8 3x -y =4 3x - y =4 What are they?

OpenStudy (anonymous):

In fact the second equation is exactly the same as the first after simplification. Hence they are dependent and consistent.